# Identifying Causal Direction via Variational Bayesian Compression

### v3

@@ Appendix B Prior and Variational Distributions
$p\left(\mathbf{b}\right)$ $=\prod_{j}p\left(b_{j}\right),$ $p\left(b_{j}\right)$ $=\mathcal{N}\left(b_{j}\mid 0,1\right).$ (24)
@@
Let $\mathbf{z}=\left[z_{ij}\right]$ denote the scale hyperparameters of the priors of the weight matrices as in Eq. (23). The marginal likelihood can be acquired by marginalizing over $\mathbf{z}$ as
[…]
@@
where $\text{where }\hat{\mathbf{z}}$ $\coloneqq\operatorname*{\arg\max}_{\mathbf{z}}\left[\log p\left(y^{\left(1:N\right)},\mathbf{z}\mid x^{\left(1:N\right)}\right)\right],\text{ and}$ (27)
@@
We also follow Dhir et al. (2024a) by choosing a uniform hyperprior over the hyperparameters $\mathbf{z}$ and assuming a sufficient amount of samples so that the distribution concentrates around a single point. Given these conditions, we can simply approximate $\log p\left(y^{\left(1:N\right)}\mid x^{\left(1:N\right)}\right)\approx\log p\left(y^{\left(1:N\right)}\mid x^{\left(1:N\right)},\hat{\mathbf{z}}\right)$. From an implementation perspective, this means that we can regard the priors $\mathbf{z}$ as additional parameters to be optimized in conjunction with the following variational parameters.
@@
1:  $\mathbf{M}_{\text{out}}\leftarrow\mathbf{H}_{\text{in}}\mathbf{M}_{\mathbf{W}}+\boldsymbol{\mu}_{\mathbf{b}}$
2:  $\mathbf{V}_{\text{out}}\leftarrow\mathbf{H}_{\text{in}}^{\circ 2}\mathbf{V}_{\mathbf{W}}+\boldsymbol{\sigma}_{\mathbf{b}}^{2}$
3:  $\mathbf{E}\sim\mathcal{N}\left(0,1\right)$dims: $N\times B$}
4:  $\mathbf{H}_{\text{out}}\leftarrow\mathbf{M}_{\text{out}}+\mathbf{V}_{\text{out}}^{\circ\frac{1}{2}}\odot\mathbf{E}$
@@ C.2 Causal Identifiability of Marginal Likelihoods
If the two causal models are distribution-equivalent, the maximum likelihoods cannot distinguish the causal direction. Otherwise, there exists some dataset $\mathcal{D}^{N}$ ($N$ sufficiently large) such that the maximum likelihood can determine the cause direction. For Bayesian causal models, which include auxiliary prior distributions over the marginal distributions ($\mathcal{C}_{X}$ and $\mathcal{C}_{Y}$) and the conditional distributions ($\mathcal{C}_{Y\mid X}$ and $\mathcal{C}_{X\mid Y}$), we have a definition of Bayesian distribution-equivalent causal models as follows:
@@
where $\mathbf{w}_{\mathbf{h}_{Y}}=\left[w_{\mathbf{h}_{Y},i_{\mathbf{h}}}\right]\in\mathbb{R}^{1\times D_{\mathbf{h}}}$, $w_{\mathbf{h}_{Y},i_{\mathbf{h}}}$ is sampled from a Gaussian prior $\mathcal{N}\left(0,z_{\mathbf{h}_{Y},i_{\mathbf{h}}}^{2}\right)$, $\mathbf{b}_{\mathbf{h}_{Y}}=\left[b_{\mathbf{h}_{Y},i_{\mathbf{h}}}\right]\in\mathbb{R}^{D_{\mathbf{h}}}$, $b_{\mathbf{h}_{Y},i_{\mathbf{h}}}\sim\mathcal{N}\left(0,1\right)$, and $\sigma\left(\cdot\right)$ is an activation function such as the hyperbolic tangent ($\tanh$). From the features $\mathbf{h}_{Y}$ in the hidden layer, the output function $f_{Y}$ is then linearly represented as
@@
where $\mathbf{w}_{f_{Y}}=\left[w_{f_{Y},i_{\mathbf{h}}}\right]\in\mathbb{R}^{D_{\mathbf{h}}\times 1}$, $w_{f_{Y},i_{\mathbf{h}}}$ is also sampled from a Gaussian distribution $\mathcal{N}\left(0,z_{f_{Y},i_{\mathbf{h}}}^{2}\right)$, and $b_{f_{Y}}\sim\mathcal{N}\left(0,1\right)$. Let us denote that $\boldsymbol{\theta}_{f_{Y}}=\left(\mathbf{w}_{f_{Y}},b_{f_{Y}}\right)$, $\boldsymbol{\theta}_{\mathbf{h}_{Y}}=\left(\mathbf{w}_{\mathbf{h}_{Y}},\mathbf{b}_{\mathbf{h}_{Y}}\right)$, $\mathbf{z}^{2}_{f_{Y}}=\left[z^{2}_{f_{Y},i_{\mathbf{h}}}\right]$, and $\mathbf{z}^{2}_{\mathbf{h}_{Y}}=\left[z^{2}_{\mathbf{h}_{Y},i_{\mathbf{h}}}\right]$, the Gaussian process of $f_{Y}$ is $\mathcal{GP}\left(m_{f_{Y}}\left(x\right),K_{f_{Y}}\left(x,x^{\prime}\right)\right)$ where
@@
$=\mathbb{E}_{\boldsymbol{\theta}_{\mathbf{h}_{Y}}}\left[\mathbf{h}_{Y}^{\top}\operatorname{diag}\left(\mathbf{z}_{f_{Y}}^{2}\right)\mathbf{h}_{Y}^{\prime}\right]+1$
$=\mathbb{E}_{\boldsymbol{\theta}_{\mathbf{h}_{Y}}}\left[{\mathbf{h}_{Y}\left(x\right)}^{\top}\operatorname{diag}\left(\mathbf{z}_{f_{Y}}^{2}\right)\mathbf{h}_{Y}\left(x^{\prime}\right)\right]+1.$ (39)
@@
[…]
$p\left(y\mid x,\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)$ $=\frac{p\left(x\mid y,\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)\left(y\mid M_{Y\rightarrow X}\right)}{p\left(x\mid\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)}.$ (48)
@@
[…]
@@
where $K_{f_{Y,i}}\left(x,x^{\prime}\right)=\mathbb{E}_{\boldsymbol{\theta}_{\mathbf{h}_{Y}}}\left[{\mathbf{h}_{Y}\left(x\right)}^{\top}\operatorname{diag}\left(\mathbf{z}_{f_{Y,i}}^{2}\right)\mathbf{h}_{Y}\left(x^{\prime}\right)\right]+1$ and $i\in\left\{1,2\right\}$.
@@ Appendix D Implementations & Hyperparameters
All the experiments are conducted on a workstation with an Intel® Core™ i7 processor, 64 GB of memory, and 3 TB of storage, except for those involving GPLVM, which are executed on NVIDIA® Tesla® V100 GPUs. Since there are no significant differences in results when performing the experiments on the baselines and our approach repeatedly, we do not include the error bars in the reported results.
@@ D.1 COMIC
Our proposed method—COMIC—is implemented in Python with the PyTorch library (Paszke et al., 2019). The implementation of the fully-connected Bayesian layers with Gaussian priors is adapted from the publicly available code [^4] Bayesian layers: https://github.com/KarenUllrich/Tutorial_BayesianCompressionForDL of Louizos et al. (2017). As mentioned in App. C, Bayesian neural networks are employed to model the mean and the standard deviation of each assumed direction, i.e., $X\rightarrow Y$ and $Y\rightarrow X$. Each neural network includes one hidden layer with $D_{\mathbf{h}}=20$ nodes with the hyperbolic tangent ($\tanh$) as the activation function and a fully-connected output layer. The softplus function is chosen as the positive link function for ensuring the positivity of the predicted standard deviation values. The neural networks of COMIC are optimized to minimize the variational Bayesian objective with the Adam optimizer (Kingma & Ba, 2015), a learning rate of $10^{-2}$, and $T=1,000$ training epochs. To avoid the bad local optima of the variational objective, Louizos et al. (2017) suggest the adoption of the “warm-up” scheme from Sønderby et al. (2016), where the model complexity term $\mathrm{KL}\left(q_{\phi}\left(w\right)\mid\mid p_{W}\left(w\right)\right)$ is weighted by a hyperparameter $\beta$ annealed linearly from 00 to $1$ for the first $T_{WU}$ training epochs. In our implementation, we choose $T_{WU}=100$. All samples of each pair are utilized in the training process, i.e., we set the batch size to be equal to the number of available samples of each dataset.
@@ D.2 Baselines
For information theory-based methods including Sloppy [^5] Sloppy: https://eda.rg.cispa.io/prj/sloppy/ (Marx & Vreeken, 2019a), Slope [^6] Slope: https://eda.rg.cispa.io/prj/slope/ (Marx & Vreeken, 2017, 2019b), and QCCD [^7] QCCD: https://github.com/tagas/QCCD (Tagasovska et al., 2020), we utilize their original repositories in R. The rpy2 interface is employed to incorporate these methods into our source code. As a Python version of IGCI (Daniušis et al., 2010; Janzing et al., 2012) is available in the Causal Discovery Toolkit (CDT [^8] CDT: https://github.com/FenTechSolutions/CausalDiscoveryToolboxKalainathan et al., 2020), we employ this version for our experiments. With GPLVM [^9] GPLVM: https://github.com/Anish144/causal_discovery_bayesian_model_selection, we utilize the Python implementation from the repository of Dhir et al. (2024a). Following to the recommendation of Dhir et al. (2024a), we choose the closed form GPLVM for all datasets except for the Tübingen dataset where the stochastic GPLVM is employed.
For regression-based methods, the Python source code of LOCI [^10] LOCI: https://github.com/aleximmer/loci is obtainable from the repository of Immer et al. (2023). The R version of CAM [^11] CAM: https://github.com/cran/CAM (Bühlmann et al., 2014) can be accessed via the Comprehensive R Archive Network (CRAN). RESIT [^12] RESIT: https://staff.fnwi.uva.nl/j.m.mooij/code/codeANM.zip is implemented in R by Peters et al. (2014), which is available on the website of the author. Both CAM and RESIT are evaluated with generalized additive models (GAMs) as their regression models.
@@ Appendix E Descriptions of the Benchmarks
All the benchmarks are available in the repository of LOCI (Immer et al., 2023) [^13] See Fn. 10.
@@ Appendix F Detailed Experimental Results
Table 1: Detailed accuracy and bidirectional AUROC scores of our COMIC approach and baseline methods across all benchmarks. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. The highest scores of each benchmark are listed in bold. Our COMIC method is compared against Sloppy (including the AIC and BIC variants, Marx & Vreeken, 2019a), Sloper (Marx & Vreeken, 2019b) and Slope (Marx & Vreeken, 2017), QCCD (Tagasovska et al., 2020), IGCI (with uniform and Gaussian reference measures, Daniušis et al., 2010), GPLVM (Dhir et al., 2024a), LOCI (with HSIC and maximum likelihood, Immer et al., 2023), CAM (Bühlmann et al., 2014), and RESIT (Peters et al., 2014). Our method outperforms the baselines on most benchmarks, except for CE-Multi.
Method AN AN-s LS LS-s MN-U SIM SIM-c SIM-G SIM-ln CE-Multi CE-Net CE-Cha Tübingen
@@
COMIC (Ours) $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{0.90}$ $\mathbf{0.93}$ $\mathbf{0.99}$ $\mathbf{1.00}$ $0.89$ $\mathbf{0.97}$ $\mathbf{0.90}$ $\mathbf{0.87}$
Sloppy-AIC $0.75$ $0.31$ $0.71$ $0.12$ $0.04$ $0.54$ $0.64$ $0.52$ $0.57$ $0.90$ $0.69$ $0.58$ $0.71$
Sloppy-BIC $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.56$ $0.96$ $0.64$ $0.62$ $0.81$ $0.77$ $0.46$ $0.79$ $0.49$ $0.61$
Sloper $0.91$ $0.56$ $0.80$ $0.06$ $0.10$ $0.57$ $0.66$ $0.60$ $0.81$ $0.86$ $0.73$ $0.59$ $0.67$
Slope $0.18$ $0.28$ $0.20$ $0.12$ $0.07$ $0.46$ $0.54$ $0.46$ $0.47$ $0.89$ $0.62$ $0.56$ $0.72$
QCCD $\mathbf{1.00}$ $0.85$ $\mathbf{1.00}$ $0.99$ $0.99$ $0.64$ $0.72$ $0.66$ $0.85$ $0.52$ $0.82$ $0.54$ $0.77$
IGCI-Uniform $0.20$ $0.35$ $0.46$ $0.34$ $0.11$ $0.37$ $0.45$ $0.53$ $0.51$ $0.92$ $0.56$ $0.55$ $0.68$
IGCI-Gaussian $0.89$ $0.97$ $0.95$ $0.94$ $0.86$ $0.36$ $0.42$ $0.86$ $0.59$ $0.68$ $0.57$ $0.55$ $0.63$
GPLVM $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.83$ $0.79$ $0.92$ $0.90$ $\mathbf{0.96}$ $0.96$ $0.75$ $0.70$
LOCI-HSIC $\mathbf{1.00}$ $\mathbf{1.00}$ $0.94$ $0.89$ $\mathbf{1.00}$ $0.78$ $0.82$ $0.78$ $0.73$ $0.79$ $0.86$ $0.73$ $0.63$
LOCI-Likelihood $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.49$ $0.50$ $0.78$ $0.79$ $0.72$ $0.77$ $0.43$ $0.56$
CAM $\mathbf{1.00}$ $\mathbf{1.00}$ $0.98$ $0.52$ $0.86$ $0.56$ $0.62$ $0.82$ $0.87$ $0.35$ $0.78$ $0.53$ $0.52$
RESIT $\mathbf{1.00}$ $\mathbf{1.00}$ $0.62$ $0.04$ $0.01$ $0.77$ $0.83$ $0.76$ $0.79$ $0.37$ $0.79$ $0.73$ $0.62$

### v4

@@ Appendix B Prior and Variational Distributions
$p\left(\mathbf{b}\right)$ $=\prod_{j}p\left(b_{j}\right),$ $p\left(b_{j}\right)$ $=\mathcal{N}\left(b_{j}\mid 0,\sigma_{b}^{2}\right).$ (24)
@@
Let $\mathbf{z}=\left[z_{ij},\sigma_{b}\right]$ denote the scale hyperparameters of the priors as in Eq. (23) and (24). The marginal likelihood can be acquired by marginalizing over $\mathbf{z}$ as
$\log p\left(y^{\left(1:N\right)}\mid x^{\left(1:N\right)}\right)=\log\int_{\mathbf{z}}p\left(y^{\left(1:N\right)}\mid x^{\left(1:N\right)},\mathbf{z}\right)p\left(\mathbf{z}\right)d\mathbf{z},$ (25)
@@
$\text{where }\hat{\mathbf{z}}$ $\coloneqq\operatorname*{\arg\max}_{\mathbf{z}}\left[\log p\left(y^{\left(1:N\right)},\mathbf{z}\mid x^{\left(1:N\right)}\right)\right],\text{ and}$ (27)
@@
[…]
@@
[…]
@@ C.2 Causal Identifiability of Marginal Likelihoods
If the two causal models are distribution-equivalent, the maximum likelihoods cannot distinguish the causal direction. Otherwise, there exists some dataset $\mathcal{D}^{N}$ (NN sufficiently large) such that the maximum likelihood can determine the cause direction. For Bayesian causal models, which include auxiliary prior distributions over the marginal distributions ($\mathcal{C}_{X}$ and $\mathcal{C}_{Y}$) and the conditional distributions ($\mathcal{C}_{Y\mid X}$ and $\mathcal{C}_{X\mid Y}$), we have a definition of Bayesian distribution-equivalent causal models as follows:
@@
where $\mathbf{w}_{\mathbf{h}_{Y}}=\left[w_{\mathbf{h}_{Y},i_{\mathbf{h}}}\right]\in\mathbb{R}^{1\times D_{\mathbf{h}}}$, $w_{\mathbf{h}_{Y},i_{\mathbf{h}}}$ is sampled from a Gaussian prior $\mathcal{N}\left(0,z_{\mathbf{h}_{Y},i_{\mathbf{h}}}^{2}\right)$, $\mathbf{b}_{\mathbf{h}_{Y}}=\left[b_{\mathbf{h}_{Y},i_{\mathbf{h}}}\right]\in\mathbb{R}^{D_{\mathbf{h}}}$, $b_{\mathbf{h}_{Y},i_{\mathbf{h}}}\sim\mathcal{N}\left(0,\sigma_{b_{\mathbf{h}}}^{2}\right)$, and $\sigma\left(\cdot\right)$ is an activation function such as the hyperbolic tangent ($\tanh$). From the features $\mathbf{h}_{Y}$ in the hidden layer, the output function $f_{Y}$ is then linearly represented as
@@
where $\mathbf{w}_{f_{Y}}=\left[w_{f_{Y},i_{\mathbf{h}}}\right]\in\mathbb{R}^{D_{\mathbf{h}}\times 1}$, $w_{f_{Y},i_{\mathbf{h}}}$ is also sampled from a Gaussian distribution $\mathcal{N}\left(0,z_{f_{Y},i_{\mathbf{h}}}^{2}\right)$, and $b_{f_{Y}}\sim\mathcal{N}\left(0,\sigma_{b_{f_{Y}}}^{2}\right)$. Let us denote that $\boldsymbol{\theta}_{f_{Y}}=\left(\mathbf{w}_{f_{Y}},b_{f_{Y}}\right)$, $\boldsymbol{\theta}_{\mathbf{h}_{Y}}=\left(\mathbf{w}_{\mathbf{h}_{Y}},\mathbf{b}_{\mathbf{h}_{Y}}\right)$, $\mathbf{z}^{2}_{f_{Y}}=\left[z^{2}_{f_{Y},i_{\mathbf{h}}},\sigma^{2}_{f_{Y}}\right]$, and $\mathbf{z}^{2}_{\mathbf{h}_{Y}}=\left[z^{2}_{\mathbf{h}_{Y},i_{\mathbf{h}}},\sigma^{2}_{\mathbf{h}_{Y}}\right]$, the Gaussian process of $f_{Y}$ is $\mathcal{GP}\left(m_{f_{Y}}\left(x\right),K_{f_{Y}}\left(x,x^{\prime}\right)\right)$ where
@@
$=\mathbb{E}_{\boldsymbol{\theta}_{\mathbf{h}_{Y}}}\left[\mathbf{h}_{Y}^{\top}\operatorname{diag}\left(\mathbf{z}_{f_{Y}}^{2}\right)\mathbf{h}_{Y}^{\prime}\right]+\sigma_{b_{f_{Y}}}^{2}$
$=\mathbb{E}_{\boldsymbol{\theta}_{\mathbf{h}_{Y}}}\left[{\mathbf{h}_{Y}\left(x\right)}^{\top}\operatorname{diag}\left(\mathbf{z}_{f_{Y}}^{2}\right)\mathbf{h}_{Y}\left(x^{\prime}\right)\right]+\sigma_{b_{f_{Y}}}^{2}.$ (39)
@@
$p\left(x\mid\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)$ $=\int p\left(x\mid y,\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)p\left(y\mid M_{Y\rightarrow X}\right)dy$ (47)
$p\left(y\mid x,\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)$ $=\frac{p\left(x\mid y,\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)p\left(y\mid M_{Y\rightarrow X}\right)}{p\left(x\mid\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)}.$ (48)
@@
$\mathbb{E}_{p\left(x\mid\mathbf{z}_{f_{X}}^{2},\mathbf{z}_{\mathbf{h}_{X}}^{2},M_{Y\rightarrow X}\right)}\left[x^{2}\right]$ $=\int\mathcal{N}\left(y\mid 0,1\right)\left(\int x^{2}\mathcal{N}\left(x\mid 0,K_{f_{X}}\left(y,y\right)+\sigma_{\varepsilon_{X}}^{2}\right)dx\right)dy$ (49)
$=\int\mathcal{N}\left(y\mid 0,1\right)\left(K_{f_{X}}\left(y,y\right)+\sigma_{\varepsilon_{X}}^{2}\right)dy.$ (50)
@@
where $K_{f_{Y,i}}\left(x,x^{\prime}\right)=\mathbb{E}_{\boldsymbol{\theta}_{\mathbf{h}_{Y}}}\left[{\mathbf{h}_{Y}\left(x\right)}^{\top}\operatorname{diag}\left(\mathbf{z}_{f_{Y,i}}^{2}\right)\mathbf{h}_{Y}\left(x^{\prime}\right)\right]+\sigma_{b_{f_{Y,i}}}^{2}$ and $i\in\left\{1,2\right\}$.
@@ Appendix D Implementations & Hyperparameters
All the experiments are conducted on a workstation with an Intel® Core™ i7 processor, 64 GB of memory, and 3 TB of storage, except for those involving GPLVM, which are executed on NVIDIA® Tesla® V100 GPUs.
@@ D.1 COMIC
Our proposed method—COMIC—is implemented in Python with the PyTorch library (Paszke et al., 2019). The implementation of the fully-connected Bayesian layers with Gaussian priors is adapted from the publicly available code [^3] Bayesian layers: https://github.com/KarenUllrich/Tutorial_BayesianCompressionForDL of Louizos et al. (2017). As mentioned in App. C, Bayesian neural networks are employed to model the mean and the standard deviation of each assumed direction, i.e., $X\rightarrow Y$ and $Y\rightarrow X$. Each neural network includes one hidden layer with $D_{\mathbf{h}}=50$ nodes with the hyperbolic tangent ($\tanh$) as the activation function and a fully-connected output layer. The natural exponential function is chosen as the positive link function for ensuring the positivity of the predicted standard deviation values. The neural networks of COMIC are optimized to minimize the variational Bayesian objective with the Adam optimizer (Kingma & Ba, 2015) and the cosine learning rate scheduler with a maximum learning rate of $10^{-2}$, a minimum learning rate of $10^{-6}$, and $T=2,500$ training epochs. To avoid the bad local optima of the variational objective, Louizos et al. (2017) suggest the adoption of the “warm-up” scheme from Sønderby et al. (2016), where the model complexity term $\mathrm{KL}\left(q_{\phi}\left(w\right)\mid\mid p_{W}\left(w\right)\right)$ is weighted by a hyperparameter $\beta$ annealed linearly from 0 to $1$ for the first $T_{\text{WU}}$ training epochs. In our implementation, we choose $T_{\text{WU}}=250$. In addition, to avoid overfitting the hyperparameters of the priors, we pretrain the models and the hyperparameters with maximum a posteriori (MAP) estimation before the variational optimization for $T_{\text{WU}}=2,500$ epochs, where the parameters and hyperparameters are optimized to minimize the negative joint log-likelihood of the data and the parameters. All samples of each pair are utilized in the training process, i.e., we set the batch size to be equal to the number of available samples of each dataset.
@@ D.2 Baselines
For information theory-based methods including Sloppy [^4] Sloppy: https://eda.rg.cispa.io/prj/sloppy/ (Marx & Vreeken, 2019a), Slope [^5] Slope: https://eda.rg.cispa.io/prj/slope/ (Marx & Vreeken, 2017, 2019b), and QCCD [^6] QCCD: https://github.com/tagas/QCCD (Tagasovska et al., 2020), we utilize their original repositories in R. The rpy2 interface is employed to incorporate these methods into our source code. As a Python version of IGCI (Daniušis et al., 2010; Janzing et al., 2012) is available in the Causal Discovery Toolkit (CDT [^7] CDT: https://github.com/FenTechSolutions/CausalDiscoveryToolboxKalainathan et al., 2020), we employ this version for our experiments. With GPLVM [^8] GPLVM: https://github.com/Anish144/causal_discovery_bayesian_model_selection, we utilize the Python implementation from the repository of Dhir et al. (2024a). Following to the recommendation of Dhir et al. (2024a), we choose the closed form GPLVM for all datasets except for the Tübingen dataset where the stochastic GPLVM is employed. For regression-based methods, the Python source code of LOCI [^9] LOCI: https://github.com/aleximmer/loci is obtainable from the repository of Immer et al. (2023). The R version of CAM [^10] CAM: https://github.com/cran/CAM (Bühlmann et al., 2014) can be accessed via the Comprehensive R Archive Network (CRAN).
@@ Appendix E Descriptions of the Benchmarks
All the benchmarks are available in the repository of LOCI (Immer et al., 2023) [^11] See Fn. 9.
@@ Appendix F Detailed Experimental Results
Table 1: Detailed accuracy and bidirectional AUROC scores of our COMIC approach and baseline methods across all benchmarks. Higher accuracy and bidirectional AUROC (Bi-AUROC) are preferable. The highest scores are listed in bold, the second-best overall results are listed in italic. Our COMIC method is compared against Sloppy (including the AIC and BIC variants, Marx & Vreeken, 2019a), Sloper (Marx & Vreeken, 2019b) and Slope (Marx & Vreeken, 2017), QCCD (Tagasovska et al., 2020), IGCI (with uniform and Gaussian reference measures, Daniušis et al., 2010), GPLVM (Dhir et al., 2024a), LOCI (Immer et al., 2023), and CAM (Bühlmann et al., 2014). Our COMIC method ranks second overall while maintaining decently consistent performance, as indicated by its lower standard deviations compared to most baseline methods.
Method AN AN-s LS LS-s MN-U SIM SIM-c SIM-G SIM-ln CE-Multi CE-Net CE-Cha Tübingen Overall
@@
COMIC (Ours) $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.57$ $0.54$ $0.78$ $0.89$ $0.79$ $0.78$ $0.43$ $0.67$ $\mathit{0.804\pm 0.200}$
Sloppy-AIC $0.75$ $0.31$ $0.71$ $0.12$ $0.04$ $0.54$ $0.64$ $0.52$ $0.57$ $0.90$ $0.69$ $0.58$ $0.71$ $0.544\pm 0.249$
Sloppy-BIC $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.56$ $0.96$ $0.64$ $0.62$ $0.81$ $0.77$ $0.46$ $0.79$ $0.49$ $0.61$ $0.749\pm 0.199$
Sloper $0.91$ $0.56$ $0.80$ $0.06$ $0.10$ $0.57$ $0.66$ $0.60$ $0.81$ $0.86$ $0.73$ $0.59$ $0.67$ $0.609\pm 0.260$
Slope $0.18$ $0.28$ $0.20$ $0.12$ $0.07$ $0.46$ $0.54$ $0.46$ $0.47$ $0.89$ $0.62$ $0.56$ $0.72$ $0.429\pm 0.246$
QCCD $\mathbf{1.00}$ $0.85$ $\mathbf{1.00}$ $0.99$ $0.99$ $0.64$ $0.72$ $0.66$ $0.85$ $0.52$ $0.82$ $0.54$ $\mathbf{0.77}$ $0.796\pm 0.171$
IGCI-Uniform $0.20$ $0.35$ $0.46$ $0.34$ $0.11$ $0.37$ $0.45$ $0.53$ $0.51$ $0.92$ $0.56$ $0.55$ $0.68$ $0.464\pm 0.207$
IGCI-Gaussian $0.89$ $0.97$ $0.95$ $0.94$ $0.86$ $0.36$ $0.42$ $0.86$ $0.59$ $0.68$ $0.57$ $0.55$ $0.63$ $0.712\pm 0.210$
GPLVM $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{0.83}$ $0.79$ $\mathbf{0.92}$ $\mathbf{0.90}$ $\mathbf{0.96}$ $\mathbf{0.96}$ $\mathbf{0.75}$ $0.70$ $\mathbf{0.909\pm 0.105}$
LOCI $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{1.00}$ $0.49$ $0.50$ $0.78$ $0.79$ $0.72$ $0.77$ $0.43$ $0.56$ $0.773\pm 0.219$
CAM $\mathbf{1.00}$ $\mathbf{1.00}$ $0.98$ $0.52$ $0.86$ $0.56$ $0.62$ $0.82$ $0.87$ $0.35$ $0.78$ $0.53$ $0.52$ $0.725\pm 0.217$